Same as above, the standard deviation of the sample means is equal to: a) 2.45 22. What is the effect of choosing a larger sample size for the sampling distribution? a) mean increases, standard deviation unchanged 23.
Answer #1

23)

Let we have samples X_{1},X_{2},...,X_{n} from the distribution having mean \mu and standard deviation \sigma then according to sampling distribution of sample mean \overline{x} is \overline{x}\sim AN(\mu_{\overline{x}}=\mu,\sigma_{\overline{x}}=\sigma/\sqrt{n})

So if we choose large sample size n i.e. if we increase the sample size n mean \mu_{\overline{x}} is unchanged and SD is decrease i.e if n\rightarrow \infty then \sigma/\sqrt{n}\rightarrow 0

hence mean unchanged, standard deviation decreases

Note Q22 requires additional data (sd . n)

sd(sample mean) = sd/sqrt(n)  

Please rate